OpenStudy (anonymous):

How many grams of tungsten trioxide must you start with to prepare 6.60g of tungsten? (For WO3, MW = 231.8 amu.)

OpenStudy (zpupster):

Write a balanced equation: \[2WO_3 \rightarrow 2W + 3O_2\] find moles of W mwt 183.8 g/mole \[\left( 6.6g W \right)\left( 1 Mole/183.8 g \right) = .036 moles\] The balanced equation tells us that 2 moles of W come from 2 moles of WO3. so we need .036 moles of WO3 \[(.036 moles of WO_3)\left( 231.8gof WO_3/mol \right)=8.35 g of WO_3\]